OpenStudy (anonymous):

Solve for x in the proportion 9 over 27 times x equals x over the quantity x plus 3

OpenStudy (anonymous):

x = 2 and x = 0 x ≈ −2.48 and x ≈ 2.48 x ≈ −0.85 and x ≈ 1.18 x ≈ −6 and x ≈ 6.16

OpenStudy (anonymous):

@jhonyy9 ?(:

jhonyy9 (jhonyy9):

9 x --- = ----- 27x x+3 hence is correct ?

jhonyy9 (jhonyy9):

not is hope wan being 9x/27 = x/(x+3) x/3 = x/(x+3) x(x+3)= 3x x^2 +3x = 3x x^2 =0 x =0 how is writing correct ?